OpenStudy (anonymous):

You are building a rectangular planter for your school garden. You want the area of the bottom to be 90ft^2. You want the length of the planter to be 3 ft longer than twice its width. What should the dimensions of the bottom of the planter be?

OpenStudy (anonymous):

are there options?

OpenStudy (anonymous):

No, there are not answer choices :(

OpenStudy (amorfide):

you know that the area of a rectangle is LW where L is length W is width so LW=90 you know the length is 3 ft longer than 2 x the width so L=2W+3 you get an equation of two variables LW=90 L=2W+3 substitute L into LW=90 to get (2W+3)W=90

OpenStudy (amorfide):

work out W and L

OpenStudy (anonymous):

What do you mean by work out W and L?? @amorfide

OpenStudy (amorfide):

you have an equation of W (2W+3)W=90 work out the value for W then work out the value for L and that is what the dimensions should be L and W